2. Table 2 presents the representative income elasticities for three categories of goods: Transportation,
    Food, and nonfed Ground Beef. (Note: Nonfed ground beef comes from cattle that have not been fed a special
    diet to produce more tender beef. Most cattle are fed corn for 90 to 120 days before going to market and thus
    produce more tender beef than nonfed cattle.)

In Table 2 below, the elasticity is interpreted as the percentage change in the quantity demanded of the good in response to a 1% increase in consumer income.  


Table 2

Goods

Income Elasticity

Transportation

1.80

Food

0.80

Ground Beef, nonfed

-1.94

a. Interpret the elasticity information in Table 2 to identify whether each of the following categories of goods is a normal or inferior good:

1) Transportation

2) Food

3) Ground Beef, non-fed

b. Your research department has estimated the income elasticity of demand for ground beef as listed in Table 2
    above.Due to an upturn in the economy, consumer incomes are expected to rise by 10 percent over the next three years. As a manager of a meat-processing plant, how will this affect your purchases of non-fed cattle? Compute the percentage change in the quantity demanded for non-fed ground beef.   

c. The inventory management team at a grocery store chain has estimated the income elasticity of demand for food
    as listed in Table 2. Due to a downturn in the economy, consumer incomes are expected to fall by 3% over the next year. As the inventory control manager for the chain, how will this affect your purchases of food? Compute the percentage change in the quantity demanded of food in response to the change in consumer incomes.
Answer #1

a. 1) Normal
2) Normal
3) Inferior
(Good is normal when income elasticity is positive and inferior when income elasticity is inferior)

b. Income elasticity of demand for ground beef = Percentage change in quantity demanded for non-fed ground beef/Percentage change in income = -1.94
So, Percentage change in quantity demanded for non-fed ground beef = (-1.94)*(Percentage change in income) = (-1.94)*(10%) = -19.4%
So, quantity demanded of ground beef will decrease by 19.4%

c. Income elasticity of demand for food = Percentage change in quantity demanded for food/Percentage change in income = 0.80
So, Percentage change in quantity demanded for food = (0.80)*(Percentage change in income) = (0.80)*(-3%) = -2.4%
So, quantity demanded of food will decrease by 2.4%